# Steerability of Instrumental-Convergence Tendencies in LLMs

### v1

@@ 5 Limitations
InstrumentalEval is small (76 scenarios), the judge evaluation is automatic, and the measured rates are sensitive to non-conceptual prompt wording and other evaluation choices. Accordingly, we avoid over-interpreting small changes: differences of a few percentage points should be read as “no clear shift under this protocol” rather than evidence for or against any underlying tendency. The main signal we emphasize is the presence (or absence) of large, robust shifts (e.g., gaps on the order of tens of percentage points). In particular, strong prompt steerability in current models does not imply that steerability will remain high for future, more capable or more agentic systems, nor does it rule out control-collapse mechanisms. The main takeaway is about the present and what can be empirically analyzed.
Moreover, presented experiments use the simplest steering technique (manual prompt suffixes) to test whether instrumental-like behaviors are prompt-elicitable and prompt-suppressible [arxiv2509_manipulating, universal_steering2025] . We expect that more sophisticated steering methods (e.g., additional post-training such as SFT/RL, or representation-level interventions such as activation steering / concept editing) [arxiv2410_conceptors, arxiv2501_fgaa] could further increase or decrease measured convergence rates, and we leave systematic evaluation of such methods to future work.
@@ 6 Conclusion
We treat steerability as an empirical property: how strongly feasible interventions can shift model behavior under a stated elicitation and scoring setup. We emphasize that steerability is not a single axis. It splits into authorized steerability (benign operators reliably enforcing intended behavior) and unauthorized steerability (attackers eliciting disallowed behavior). This split makes a central tension for open-weight systems explicit: improving controllability for builders can also lower the cost of misuse.
Using InstrumentalEval, we show that minimal inference-time steering (prompt suffixes) can strongly suppress or amplify outputs labeled as instrumental convergence under our protocol. In this regime, increased model capability does not imply weaker control; aligned variants remain highly responsive to constraint prompts, and scaling can coincides with lower convergence. Consequently, for open-weight models, preventing unauthorized steering remains a major open technical problem. [security_steerability2025, arxiv2412_durability, arxiv2310_managing]

### v2

@@ 5 Limitations
InstrumentalEval is small (76 scenarios), the judge evaluation is automatic, and the measured rates might be sensitive to non-conceptual prompt wording. Accordingly, we avoid over-interpreting small changes in results: differences of a few percentage points should be read as “no clear shift under this protocol” rather than strong evidence for or against the underlying tendency. The main signal we emphasize is the presence (or absence) of large, robust shifts (e.g., gaps on the order of tens of percentage points).
Moreover, presented experiments use the simplest steering technique to test whether instrumental-like behaviors are prompt-elicitable and prompt-suppressible [arxiv2509_manipulating, universal_steering2025]. We expect that more sophisticated steering methods (e.g., additional post-training such as SFT/RL, or representation-level interventions such as activation steering / concept editing) [arxiv2410_conceptors, arxiv2501_fgaa] could further increase or decrease measured convergence rates, and we leave systematic evaluation of such methods to future work.
@@ 6 Conclusion
We treat steerability as an empirical property: how strongly feasible interventions can shift model behavior under a stated elicitation and scoring setup. We emphasize that steerability is not a single axis. It splits into authorized steerability (benign operators reliably enforcing intended behavior) and unauthorized steerability (attackers eliciting disallowed behavior). Separating authorized from unauthorized steerability highlights a core tension for open-weight systems: the same mechanisms that enable reliable control by builders can also enable misuse by adversaries.
Using InstrumentalEval, we show that minimal inference-time steering (prompt suffixes) can strongly suppress or amplify outputs labeled as instrumental convergence under our protocol. In this regime, increased model capability does not imply weaker control; aligned variants remain highly responsive to constraint prompts, and scaling can coincide with lower convergence. Consequently, for open-weight models, preventing unauthorized steering remains a major open technical problem [security_steerability2025, arxiv2412_durability, arxiv2310_managing].